The length of a rectangle is 8 in. more than its width. The perimeter of the rectangle is 24 in. What are the width and length of the rectangle?
Answer by Maths68(1474)

Perimeter = p=24 inches
Length = l=w+8
Width = w=?
Length = l=?
2 length + 2 width = Perimeter of rectangle
2l+2w=p
2(w+8)+2w=24
2w+16+2w=24
4w=24-16
4w=8
w=8/4
w=2 inches

Width = w= 2 inches
Length = l= w+8= 2+8=10 inches